OpenStudy (anonymous):

does anyone know a good website with the rules for working with brackets that are powered to fractions? i need those special rules that tell you when to do what and when to not

OpenStudy (imstuck):

Do you have a problem as an example?

OpenStudy (anonymous):

yes, i have 2 functions \[2L + K=C\] and \[\sqrt{L}+\sqrt{K}=y\] i have to determine the cost function the answer in the book is \[y= 3^{1/2} \] but i get \[y=3L ^{1/4}\] from \[L ^{0.5} +(2L)^{0.5}\] and i still dont understand why sometimes the constant is multiplied by the power and sometimes not even if the constant is inside a bracket

OpenStudy (anonymous):

the answer in the book is \[y=3L ^{1/2},\] my bad!